OpenStudy (anonymous):

Can someone please explain to me how r(t)=(1-t)<1,3,-2>+t<2,-1,3> became r(t)=<1+t,3-4t,-2+5t>

OpenStudy (anonymous):

@Azteck can you lease help?!

OpenStudy (anonymous):

what are the these "<" signs doing there?

OpenStudy (anonymous):

theyre supposed to represent vectors

OpenStudy (anonymous):

the original question asks find a vectr equation and parametric equations for th line segment that joins the point P(1,3,-2) and Q(2,-1,3)

OpenStudy (anonymous):

I'm not understanding the vector parts. Sorry about this, but unfortunately, I'm not an expert at these sorts of questions. Maybe you should ask someone else. @whpalmer4 are you good at these questions at all?

OpenStudy (anonymous):

i figured it out thanks sooo much though!

OpenStudy (anonymous):

hahah I didn't even help. Well done ranyai12. Excellent work from you. (Y)

OpenStudy (anonymous):

lol but you attempted to help which is more than enough! thanks again!

OpenStudy (whpalmer4):

\[r(t) = (1-t)<1,3,-2> + t<2, -1, 3>\]You essentially treat it like distributing and collecting terms: \[r(t) = <1(1-t),3(1-t),-2(1-t)> + <2t,-t,3t> = <1-t+2t, 3-3t-t,-2+2t+3t> =\]\[= <1+t,3-4t,-2+5t>\]

OpenStudy (anonymous):

yea i realized that thanks so much though! i at first thought it had to del with matrices which is why i got confused

OpenStudy (anonymous):

deal*